OpenStudy (anonymous):

Limit of a function with direct subsitution??? Find the limit of the function by using direct substitution. lim(x^2+8x-2) lim ->2

OpenStudy (anonymous):

With direct substitution, you have \[\lim_{x\to c}f(x)=f(c)\] So in this case, \[\lim_{x\to 2}(x^2+8x-2)=(2^2)+8(2)-2\]